OpenStudy (cwtan):

\(\Huge \frac {d (\frac {4^x}{1+x^3})}{dx}\)=?

OpenStudy (anonymous):

use the quotient rule (4^x) ' = ln4 *4^x

OpenStudy (anonymous):

let \[\Large y=\frac{4^x}{1+x^3}\] take natural log of both sides \[\Large \ln (y)=\ln(\frac{4^x}{1+x^3})\] or \[\Large \ln(y)=\ln(4^x)-\ln(1+x^3)\] \[\Large \ln(y)=x*\ln(4)-\ln(1+x^3)\] just differentiate both sides with respect to x

OpenStudy (cwtan):

what i've done is \(\Huge \frac {(4^x)[(1+x^3)ln 4 -3x^2]}{(1+x^3)^2}\) can it be more simplified? or I've wrong?

OpenStudy (anonymous):

nope that's fine.

OpenStudy (cwtan):

ok thanks.